The English alphabet contains 21 consonants and fivevowels. How many strings of six lowercase letters of theEnglish alphabet contain:

a) exactly two vowels?

b) at least two vowels?


1
Expert's answer
2021-12-17T10:27:53-0500

a)

The location of the vowels can be chosen in "\dbinom{6}{2}" ways.

For each choice of locations, the locations can be filled with vowels in "5^2"   ways.

Four empty locations can be filled with consonants in "21^4" ways.


"\dbinom{6}{2}(5)^2(21)^4=72930375"

b)

The total number of strings is "26^6=308915776"

The number of strings which contain exactly vowels is "21^6=85766121"

Exactly 1vowel: The location of the vowel can be chosen in "\dbinom{6}{1}" ways.

For each choice of location, the location can be filled with a vowel in "5^1" ways.

That leaves "5" empty locations, which can be filled with consonants in "21^5"ways


"\dbinom{6}{1}(5)^1(21)^5=122523030"

The number of strings which contain at least two vowels is


"308915776-85766121-122523030"

"=100626625"